OpenStudy (anonymous):

calculating-the rise of a line on a distance-versus-time graph is 600m and the run is 3 minutes.what is the slope of the line

OpenStudy (anonymous):

\[\text{slope}=\dfrac{\text{rise}}{\text{run}}=\dfrac{600 \text{ m}}{3 \text{ min}}=200\dfrac{\text{m}}{\text{min}}\]